The mean per capita income is 21,053 dollars per annum with a standard deviation of 805 dollars per annum. What is the probability that the sample mean would differ from the true mean by less than 162 dollars if a sample of 71 persons is randomly selected? Round your answer to four decimal places.
Answer #1

µ =    21053                                      
σ =    805                                      
n=   71                                      
we need to calculate probability for ,

21053-162 ≤x ≤ 21053 + 162
20891   ≤ X ≤    21215                                  
X1 =    20891   ,    X2 =   21215                          
                                          
Z1 =   (X1 - µ )/(σ/√n) = (   20891   -   21053   ) / (   805   / √   71   ) =   -1.70  
Z2 =   (X2 - µ )/(σ/√n) = (   21215   -   21053   ) / (   805   / √   71   ) =   1.70  
                                          
P (   20891   < X <    21215   ) =    P (    -1.70   < Z <    1.70   )       
                                          
= P ( Z <    1.70   ) - P ( Z <   -1.70   ) =    0.9550   -    0.0450   =    0.9101       (answer)
excel formula for probability from z score is =NORMSDIST(Z)